OpenStudy (anonymous):

Please help solve these!!??!?!(2 Q's) solve: 10/3x + 4/3=7+x/2x solve: x+12/x+4=x/x+8

OpenStudy (mertsj):

\[\frac{10}{3x}+\frac{4}{3}=\frac{7+x}{2x}\]

OpenStudy (mertsj):

Is that the problem?

OpenStudy (anonymous):

they are fractions

OpenStudy (radar):

Did Mertsj copy the problem correctly, she asked if it was correct.

OpenStudy (anonymous):

yess

OpenStudy (mertsj):

Then multiply each term by the common denominator which is 6x

OpenStudy (radar):

Put terms containing x on the left and numerical value on the right of equal sign. 10/3x -(7+x)/2x =-4/3, Add the two fractions containing x on the left. (-1-3x)/6x = -4/3 Now cross multiply getting 24x=-3-9x 33x=-3 x=-1/11